# PROPHET: An Inferable Future Forecasting Benchmark with Causal Intervened Likelihood Estimation

### v1

@@ A.2 Construction Details
A.3 Future Forecasting Analysis Assisted by CIL
We calculate the distribution of the CIL  metric and the number of news articles over time. We regard the time span between the oldest news and the resolved date as the whole progress of a question. Then we compute the progress of each news by the percentage of its date in this progress. The results are in Figure 5. We explore some key properties of future forecasting based on these studies.
① As the approaching the resolved date, both news of high and low CIL increase. News of high CIL increase is consistent with human intuition. As time progresses, the prediction of future events will become more certain. However, we also find low CIL news increases indicating that as time progresses, there will also be an increase in the generation of misleading information. It challenges the model to resist this bias for precise predicting.
② We mainly discovery two volume distributions of news articles. The first type of distribution is characterized by a very low number of news articles early on, with a sudden surge close to the end time. The second type of distribution is characterized by a uniform distribution of news over time. This reflects two ways in which people pay attention to events. However, the first type brings difficulties for early prediction since it lack valid information at an early date.
Figure 5: In-depth analysis. The horizontal axis represents the entire prediction process.
Figure 6: Case studies.
A.4 Evaluated Methods
We introduce the methods that we evaluate in this work. For the retrieval methods:
Naive RAG: Since the news articles are long, we first summarize the news articles in advance. This RAG method then retrieves relevant news articles via embedding similarity between the question and news summary. We use all-MiniLM-L6-v2 models in SentenceTransformer [^6] https://sbert.net. After retrieving the news, we use the scratchpad prompt for reasoning.
APP: This is the method introduced by Halawi et al. (2024). It also first summarizes the news articles. Then it uses LLM to compute the relevance score. After that, it also uses scratchpad prompt for reasoning.
Rankllama: This is a retrieval method where it can understand the complicated retrieval instructions Ma et al. (2024). It uses the model to encode the question and the news articles. We use summaries of the news. After retrieval, it answers in scratchpad prompt as well.
HyDE: Given a query, this method uses an instruction-following language model (e.g., InstructGPT) to generate a "hypothetical document" that captures relevance patterns Gao et al. (2023). In event prediction scenario, we generate potential future events that could effect the answer. Then retrieve relevant news articles.
The reasoning methods are:
ScrathPAD: This is the zero-shot ScrathPAD prompting method based on LLMs. We use the scratchpad prompt introduced by Halawi et al. (2024).
CoT: Chain of Thought is a technique that enables AI models to mimic human-like step-by-step reasoning by breaking down complex problems into intermediate logical steps, significantly improving interpretability and accuracy in tasks such as mathematical reasoning and NLP Wei et al. (2022).
Long-CoT: Long-CoT is on LLMs trained with reinforcement learning to perform advanced reasoning through internal CoT such as OpenAI-O1 OpenAI (2024), achieving state-of-the-art performance in competitive programming, mathematics, and scientific benchmarks, even surpassing human experts in some domains.
Type Benchmark W G R I
Script Learning MCNC Granroth-Wilding (2016) -
SCT Mostafazadeh et al. (2017) -
CoScript Yuan et al. (2023) -
Time Series GDELT Leetaru and Schrodt (2013) -
ICEWS Schrodt et al. (2012) -
Event Reasoning ECARE Du et al. (2022) -
EV2 Tao et al. (2024) -
Open Event Prediction Halawi et al. (2024)
OpenEP Guan et al. (2024)
ForecastBench Karger et al. (2024)
PROPHET (Ours)
Table 4: Comparison with other forecasting benchmarks. W: real-world questions. G: News Grounded. R: reproductive. I: inferable validation.
Model Version
GPT-4o gpt-4o-2024-08-06
GPT-4o-mini gpt-4o-mini-2024-07-18
O1-mini o1-mini-2024-09-12
Claude claude-3-5-sonnet-20240620
Gemini gemini-1.5-pro-latest
Qwen2.5-32B Qwen2.5-32B-Instruct-GPTQ-Int4
Qwen2.5-7B Qwen2.5-7B-Instruct-GPTQ-Int4
Table 5: Evaluated model versions.
A.5 Evaluation Details
All experiments in this work are under twice runs. We report the mean and std values. We list the versions of LLMs we use in Table 5. The reasoning prompts are in prompts (g-h).
A.6 Case of CIL
In this section, we showcase articles of high and low CIL scores. In Figure 6 we illustrate two questions. Each question is paired with CIL-High and CIL-Low articles. We find our CIL estimation precisely captures supportiveness for answering the question. For example, the first question asks the CDC’s reaction to mpox. The CIL-High states the situation of vaccination of U.S. while the CIL-Low only mentions the global situation of mpox. Owing to the low vaccination rates of the U.S., it is likely that the CDC would pose the assessment of mpox exceeding "Very Low". In the second example, the CIL-High tells that Walz struggled to counter J.D. Vance effectively while CIL-Low merely mentions Kamala Harris wants to raise a debate. CIL-High contributes more to the correct answer.
A.7 Prompts
@@
(a) Entity Query Generation I will provide you with a forecasting question and the background information for the question. Extract the named entities, events of the question. Each entity and event are up to 5 words. The named entities can only be people, organizations, countries, locations while can not be date or time. Put all result items in a list that I can parse by JSON as ["entity 1", "entity 2", "event 1", "event 2", …]. Question: ${\mathcal{Q}}$ Question Background: ${\mathcal{B}}$ Question Date: date Output:
@@
(b) Resolving Steps Query Generation I will provide you with a forecasting question and the background information for the question. I will then ask you to generate short search queries (up to max words words each) that I’ll use to find articles on Google News to help answer the question. The articles should be mainly about event arguments such as subjects, objects, locations, organizations of the events in question and background information. You must generate this exact amount of queries: num keywords. Put all result items in a list that I can parse by JSON as ["step 1", "step 2", "step 3", …]. Question: ${\mathcal{Q}}$ Question Background: ${\mathcal{B}}$ Question Date: date Output:
@@
(c) Similar Events Query Generation I will provide you with a forecasting question and the background information for the question. I will then ask you to generate short search queries (up to max words words each) that I’ll use to find articles of similar events on Google News to help answer the question. The similar events are events happened on other similar entities in the history. Or events happended on question entities but on other date. You must generate this exact amount of queries: num keywords. Put all result items in a list that I can parse by JSON as ["event 1", "event 2", "event 3", …]. Question: ${\mathcal{Q}}$ Question Background: ${\mathcal{B}}$ Question Date: date Output:
(d) News Article Relevance Rating Please consider the following forecasting question and its background information. After that, I will give you a news article and ask you to rate its relevance with respect to the forecasting question. Question: ${\mathcal{Q}}$ Question Background: ${\mathcal{B}}$ Resolution Criteria: ${\mathcal{R}}$ Article: articles Please rate the relevance of the article to the question, at the scale of 1-6 1 – irrelevant 2 – slightly relevant 3 – somewhat relevant 4 – relevant 5 – highly relevant 6 – most relevant Guidelines: - If the article has events of similar types which may happened on different subjects, it also consider relevant to the question. - You don’t need to access any external sources. Just consider the information provided. - If the text content is an error message about JavaScript, paywall, cookies or other technical issues, output a score of 1. Your response should look like the following: Thoughts: { insert your thinking } Rating: { insert your rating }
(e) Conditional Probability ### Given a background that in the meantime: — These events happened: news of ${\mathcal{X}}_{n_{1}}$ — These events didn’t happen: news of ${\mathcal{X}}_{n_{2}}$## Most importantly: — These events happened: news of ${\mathcal{X}}_{i}$## Answer the question: ${\mathcal{Q}}$## Instructions: 1. Provide at least 3 reasons why the answer might be no. { Insert your thoughts } 2. Provide at least 3 reasons why the answer might be yes. { Insert your thoughts } 3. Rate the strength of each of the reasons given in the last two responses. Think like a superforecaster (e.g. Nate Silver). { Insert your rating of the strength of each reason } 4. Aggregate your considerations. { Insert your aggregated considerations } 5. Output your answer (a number between 0 and 1) with an asterisk at the beginning and end of the decimal. { Insert your answer }"
(f) Probability ### Given a situation that in the meantime: — These events happened: news of ${\mathcal{X}}_{n_{1}}$ — These events didn’t happen: news of ${\mathcal{X}}_{n_{2}}$## Instructions: Use your world knowledge and commonsense to reason the probability if the situation can happen. Generate the thoughts first: { Insert your thoughts } Then output your answer (a probability number between 0 and 1) with an asterisk at the beginning and end of the decimal. { Insert your answer }
(g) ScratchPAD Question: ${\mathcal{Q}}$ Question Background: ${\mathcal{B}}$ Resolution Criteria: ${\mathcal{R}}$ We have retrieved the following information for this question: retrieved articles Instructions: 1. Provide at least 3 reasons why the answer might be no. { Insert your thoughts } 2. Provide at least 3 reasons why the answer might be yes. { Insert your thoughts } 3. Rate the strength of each of the reasons given in the last two responses. Think like a superforecaster (e.g. Nate Silver). { Insert your rating of the strength of each reason } 4. Aggregate your considerations. { Insert your aggregated considerations } 5. Output your answer (a number between 0 and 1) with an asterisk at the beginning and end of the decimal. { Insert your answer }
(h) CoT and Long-CoT Question: ${\mathcal{Q}}$ Question Background: ${\mathcal{B}}$ Resolution Criteria: ${\mathcal{R}}$ We have retrieved the following information for this question: retrieved articles Think step by step. Reason and finally output your answer (a number between 0 and 1) with an asterisk at the beginning and end of the decimal.,

### v2

@@ 6. Conclusion
Prompts
@@
Agentic RAG Tool
We show the article retrieval tool for all agentic RAG methods:
1class AgenticRAGTool:
2 name = "query_rag"
3 description = (
4 "To search in a previously built RAG index to find the most relevant chunks of text."
5 )
6 parameters = [
7 {
8 ’name’: ’query’,
9 ’type’: ’string’,
10 ’description’: ’the query text to search for relevant text chunks.’,
11 ’required’: True
12 },
13 {
14 ’name’: ’top_k’,
15 ’type’: ’integer’,
16 ’description’: ’the number of top relevant chunks to retrieve (default is 3).’,
17 ’required’: False
18 }
19 ]
Listing 1: Python Class Definition for Agentic RAG Tool
@@
(a) Entity Query Generation I will provide you with a forecasting question and the background information for the question. Extract the named entities, events of the question. Each entity and event are up to 5 words. The named entities can only be people, organizations, countries, locations while can not be date or time. Put all result items in a list that I can parse by JSON as [”entity 1”, ”entity 2”, ”event 1”, ”event 2”, …].
Question: $Q$
Question Background: $B$
Question Date: $D$
Output:
@@
(b) Resolving Steps Query Generation I will provide you with a forecasting question and the background information for the question. I will then ask you to generate short search queries (up to max words words each) that I’$l$ use to find articles on Google News to help answer the question. The articles should be mainly about event arguments such as subjects, objects, locations, organizations of the events in question and background information. You must generate this exact amount of queries: num keywords. Put all result items in a list that I can parse by JSON as [”step 1”, ”step 2”, ”step 3”, …].
Question: $Q$
Question Background: $B$
Question Date: $D$
Output:
@@
(c) Similar Events Query Generation I will provide you with a forecasting question and the background information for the question. I will then ask you to generate short search queries (up to max words words each) that I’$l$ use to find articles of similar events on Google News to help answer the question. The similar events are events happened on other similar entities in the history. Or events happended on question entities but on other date. You must generate this exact amount of queries: num keywords. Put all result items in a list that I can parse by JSON as [”event 1”, ”event 2”, ”event 3”, …].
Question: $Q$
Question Background: $B$
Question Date: $D$
Output:
(d) News Article Relevance Rating Please consider the following forecasting question and its background information. After that, I will give you a news article and ask you to rate its relevance with respect to the forecasting question.
### Question: $Q$
### Question Background: $B$
### Article: {article}
Please rate the relevance of the article to the question, at the scale of 1-6
1 – completely irrelevant
2 – slightly relevant
3 – somewhat relevant
4 – relevant
5 – highly relevant
6 – completely relevant directly reveals the answer
Guidelines:
- scale 1-5 don’t reveal the answer, only scale 6 reveals the answer.
- If the article has events of similar types which may happened on different subjects, it also consider relevant to the question to some degree.
- You don’t need to access any external sources. Just consider the information provided.
- If the text content is an error message about JavaScript, paywall, cookies or other technical issues, output a score of 1.
Your response should look like the following, don’t generate other words in the response:
Rating:
(e) Probability $P(Y=\hat{Y}|X_{i}=1,\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})$ ### Given some event: {events}
### Most important event: {current event}
### Instructions: Based on the information, especially the most important event, compute the probability that the answer for the following question is yes. Question: $Q$
### The mentioned named entities are real in life, no need to check that. You compute the probability based on your world knowledge about the mentioned named entities and the internal logic in the situation. the Output your answer (a probability number between 0 and 1).
(f) Probability $P(\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})$ ### Given a situation describing some events: {events}
### Instructions: Reason the probability of the situation. The mentioned named entities are real in life, no need to check that. You compute the probability based on your world knowledge about the mentioned named entities and the internal logic in the situation. Output your answer (a probability number between 0 and 1).
(g) Naive RAG Your task is to answer the prediction question. The question asks about whether a future event would happen or not. Output a probability between 0 and 1 representing the likelihood of the event happening. Output the probability in the following format:
¡Probability¿…¡/Probability¿
Use the provided articles as context to inform your prediction. The articles are provided below.
Question: $Q$
Background: $B$
Create date of the question: $D$
Context Articles: {retrieved articles}
(h) Agentic RAG You are a prophet agent who can answer a query based on the data and the background information. Use the tools below to help you answer the question.
Your task is to answer the prediction question. The question asks about whether a future event would happen or not. Output a probability between 0 and 1 representing the likelihood of the event happening. Output the probability in the following format:
¡Probability¿…¡/Probability¿
Question: $Q$
Background: $B$
Create date of the question: $D$